If the sum of n terms of an A.P. is n2, then its nth term is
(a) 2n – 1 (b) 2n + 1 (c) n2 – 1 (d) 2n – 3

Grade:12

1 Answers

Vikas TU
3 years ago
Dear student
 sₙ =n/2{2a+(n-1)d}
where a =first term ,d =common difference ,n =no of terms
 sₙ=n/2(a+l)
l =last term
 
sₙ=n²
putting n= 1 , s₁=1²=1
s₁ is sum of first term which is equal to first term
a₁=s₁ =1
now putting n=2 in sₙ
s₂ =2²=4
s₂ is sum of first two terms so
a₁+a₂=s₂=4
1 + a₂=4
a₂=4-1=3
now a₁=1,d=a₂-a₁=3-1=2,n=n,aₙ=?
we have formula of nth term of AP
aₙ=a+(n-1)d
=1+(n-1)2
=1+2n-2
=(2n-1)